# PD-TPE: Parallel Decoder with Text-guided Position Encoding for 3D Visual Grounding

### v1

@@ Introduction
In order to enable the surrounding branch to focus on visual tokens containing effective layout information, we propose a text-guided method for this task to refine the attention map after position encoding. To be specific, we add a bias to $\mathcal{E}$ and update Eq. 1, which is designed as follows:
$\widehat{\mathcal{A}}=\operatorname{softmax}(\mathcal{Q}\mathcal{K}^{\text{T}}{\;+\;\mathcal{E}}{\;+\;{bias}}),$ (9)
$bias=\mathcal{F}(\operatorname{max}(V(T^{i}_{s})^{\text{T}})),i=0,1,...,L_{s},$ (10)
where $\mathcal{F}$ represents a non-linear function. $V\in\mathbb{R}^{N\times D}$ is the visual feature served as the key-value feature of visual cross-attention, $T_{s}\in\mathbb{R}^{L_{s}\times D}$ is the text surrounding feature and $L_{s}$ is the length of surrounding components from the decoupled text.
As shown in Fig. 4, we use dot product of visual and surrounding text features to compute the similarity between them, and $bias\in\mathbb{R}^{N}$ represents the maximum similarity among the $L_{s}$ text components. If a visual token is related to a specific component, it indicates that this token can provide effective surrounding layout information, so that its corresponding value in the attention map will be enhanced; otherwise, the value in the attention map will be weakened.
@@
The dense aligned loss is set up by dense position aligned loss and dense semantic aligned loss as follows:
$\mathcal{L}_{all}=\mathcal{L}_{pos}+\mathcal{L}_{sem}$ (11)
For more implementation details, please refer to (Wu et al. 2023).
@@
Unique ($\sim$19%) Multiple ($\sim$81%) Overall
Method Venue 0.25 0.5 0.25 0.5 0.25 0.5
ScanRefer (Chen, Chang, and Nießner 2020) ECCV2020 67.64 46.19 32.06 21.26 38.97 26.10
TGNN (Huang et al. 2021) AAAI2021 68.61 56.80 29.84 23.18 37.37 29.70
InstanceRefer (Yuan et al. 2021) ICCV2021 77.45 66.83 31.27 24.77 40.23 32.93
SAT (Yang et al. 2021) ICCV2021 73.21 50.83 37.64 25.16 44.54 30.14
FFL-3DOG (Feng et al. 2021) ICCV2021 78.80 67.94 35.19 25.70 41.33 34.01
3DVG-Transformer (Zhao et al. 2021) ICCV2021 77.16 58.47 38.38 28.70 45.90 34.47
3D-SPS (Luo et al. 2022) CVPR2022 84.12 66.72 40.32 29.82 48.82 36.98
3DJCG (Cai et al. 2022) CVPR2022 78.75 61.30 40.13 30.08 47.62 36.14
BUTD-DETR (Jain et al. 2022) ECCV2022 82.88 64.98 44.73 33.97 50.42 38.60
D3Net (Chen et al. 2022) ECCV2022 - 70.35 - 30.50 - 37.87
EDA (Wu et al. 2023) CVPR2023 85.76 68.57 49.13 37.64 54.59 42.26
ViewRefer (Guo et al. 2023) ICCV2023 - - 33.08 26.50 41.30 33.66
3DRP-Net (Wang et al. 2023) EMNLP2023 83.13 67.74 42.14 31.95 50.10 38.90
CORE-3DVG (Yang et al. 2023) NIPS2023 84.99 67.09 51.82 39.76 56.77 43.84
VPP-Net (Shi, Wu, and Lee 2024) CVPR2024 86.05 67.09 50.32 39.03 55.65 43.29
PD-TPE (ours) 86.75 73.01 50.12 40.62 56.84 45.65
vs. EDA (baseline) +0.99 +4.44 +0.99 +2.98 +2.25 +3.19
vs. VPP-Net +0.70 +5.92 -0.20 +1.59 +1.19 +2.36
vs. CORE-3DVG +1.76 +5.92 -1.70 +0.86 +0.07 +1.81
@@
Datasets & Metrics.
We perform experiments on two commonly used datasets, ScanRefer (Chen, Chang, and Nießner 2020) and Nr3D (Achlioptas et al. 2020). Both datasets are derived from 3D scenes within ScanNet (Dai et al. 2017), which contains 1201 indoor 3D scenes for training and 312 for validation. ScanRefer provides 51,583 human-annotated natural language descriptions of 11,046 objects in 800 3D scenes, which is officially divided into a 36,665-sample training set and a 9,508-sample testing set. Nr3D contains 41,503 language annotations, describing 76 object categories from 707 indoor scenes.
For ScanRefer, we adopt Acc@mIoU as the evaluation metric, where m is set as 0.25 and 0.5. This metric indicates the proportion of predicted bounding boxes that have an Intersection over Union (IoU) greater than m with the ground-truth bounding boxes. The results on two subsets, unique and multiple samples, are also provided, where the unique means that only one category of the target object is mentioned in the language description and multiple involves more interference candidates. For Nr3D, experiments are evaluated under the Acc@0.25IoU. The Easy and Hard subsets are divided by whether there are more than 2 distractors. The view-dependent and view-independent subsets are set based on the provided perspectives.
Implementation Details.
The code is implemented based on PyTorch. For all experiments set below, we use an AdamW optimizer with a 5e-4 weight decay and set the batch size to 12 on four 24-GB NVIDIA-4090 GPUs. For ScanRefer, we use a 2e-3 learning rate for the visual encoder and a 2e-4 learning rate for all other layers, training for 400 epochs. It takes about 15 minutes per epoch and converges at around 50 epochs. We also provide experimental results with 60 epochs to demonstrate the model’s fast convergence capability in Sec. Ablation StudiesAblation Studies. For NR3D, the learning rates are set at 1e-3 and 1e-4, 15 minutes per epoch, and 300 epochs are trained.
@@
Method Venue Easy Hard V-Dep V-Indep Overall
ReferIt3D (Achlioptas et al. 2020) ECCV2020 43.6 27.9 32.5 37.1 35.6
TGNN (Huang et al. 2021) AAAI2021 44.2 30.6 35.8 38.0 37.3
InstanceRefer (Yuan et al. 2021) ICCV2021 46.0 31.8 34.5 41.9 38.8
3DVG-Transformer (Zhao et al. 2021) ICCV2021 48.5 36.0 36.5 44.9 42.1
SAT (Yang et al. 2021) ICCV2021 56.3 42.4 46.9 50.4 49.2
MVT (Huang et al. 2022) CVPR2022 61.3 49.1 54.3 55.4 55.1
EDA (Wu et al. 2023) CVPR2023 58.2 46.1 50.2 53.1 52.1
ViewRefer (Guo et al. 2023) ICCV2023 63.0 49.7 55.1 56.8 56.0
VPP-Net (Shi, Wu, and Lee 2024) CVPR2024 - - - - 56.9
PD-TPE(Ours) 67.7 50.6 52.8 62.3 59.1
vs. EDA(Baseline) +9.5 +4.5 +2.6 +9.2 +7.0
Table 2: Performance on NR3D datasets by Acc@0.25IoU as the metric. There are some previous methods that only provide evaluation results on the det. dataset, which are not shown in this table.
Tab. 1 exhibits our experiment on the ScanRefer dataset compared with previous works. Our PD-TPE outperforms all the previous methods on Acc@0.25IoU and Acc@0.5IoU, achieving 56.84% and 45.65%, respectively. It surpasses our baseline EDA(Wu et al. 2023) by 3.19% Acc@0.5IoU, and also 1.61% higher than that of the state-of-the-art method (Yang et al. 2023). Note that compared with methods with 2D+3D inputs, our method can achieve even better results without using additional 2D features as assistance. The evaluations on both unique and multiple inputs reveal improvements, indicating that our model can effectively parse and align both complex and simple sentences.
We also report our experiment results on the Nr3d dataset in Tab. 2. All results in this table are evaluated under Acc@0.25IoU. Our PD-TPE performs better than all the previous one-stage methods and reaches 59.1%, surpasses the previous state-of-the-art method (Shi, Wu, and Lee 2024) and our baseline EDA(Wu et al. 2023) by 7.0% and 2.2%, respectively. Moreover, in the ”Hard” subset with two or more distractors that are more challenging to understand, our PD-TPE shows a remarkable performance of 50.6%, which proves that our model has strong comprehension capability when handling complex text inputs.
@@
We perform extensive ablation studies to understand the effectiveness of each component of our PD-TPE. Without further specification, all experiments are conducted on ScanRefer with 400 epochs.
Comparison of Different Relative PE Methods.
@@
In Tab. 3, we first compare the effect of different Relative Positioning methods. The experiment (a) is our baseline EDA(Wu et al. 2023). Keeping the structure of the model, especially the decoder, we add PE with different types of relative positioning methods to the decoder’s cross-attention map. The experiment (b) use Vertex Relative PE following (Shen et al. 2024), as shown in Fig. 3(a), achieving 44.64%(+2.38%) Acc@0.5IoU. As shown in experiment (c), the Center Relative PE method (Fig. 3(b)) results in limited improvement to Acc@0.5IoU (43.01%, +0.75%), because the method only considers the relative positional relationship between tokens and query proposals, without taking the shape and size of the predicted boxes into account. We finally use the Box-surface PE method in experiment (d), which reaches comparable results as experiment (b). Considering the large latency brought by Vertex Relative PE, we take the Box-surface PE method as our priority.
PE Method @0.25 @0.5 Latency/scene
(a) - 54.59 42.26 118 ms
(b) Vertex RPE 55.16 44.64 135 ms
(c) Center RPE 54.31 43.01 121 ms
(d) Box-surface RPE 54.72 44.74 123 ms
Table 3: Evaluation results and inference costs of different types of PE methods. For Latency costs per scene, we evaluate on one NVIDIA 4090 24GB GPU with batch size as 1 for a fair comparison.
Figure 5: Visualization results of different models for scenes from the ScanRefer dataset. For all boxes, red represents the ground-truth references; blue represents results from EDA that contain grounding errors; green represents proposals generated by our PD-TPE; light blue represents associated contextual objects. Words in different colors show the results of text decoupling.
Effect of the Position Encoding to the Different Structures of Decoders.
@@
We provide the ablation study on the structure of the decoder and display the results in Tab. 4. Comparing the experiment (a) and (c), our design that changes the decoder’s serial structure to a double-branch parallel structure increases the Acc@0.5IoU from 42.26% to 43.33% (+1.07%), which proves that allowing the decoder to equally focus on the features of the query proposals themselves and their surrounding layout can lead grounding more accurate. Experiments (b) and (d) add the Box-surface Relative PE to the attention map based on the decoder structure of experiments (a) and (c), respectively. The evaluation results surpass their reference by +2.48% and +1.68% Acc@0.5IoU. In experiment (e), we add a text-guided bias to the PE of the surrounding branch in the decoder. It helps the surrounding branch pay more attention to visual tokens with effective layout information, which increases the evaluation result to 45.65% Acc@0.5IoU.
Decoder Structure PE Method @0.25 @0.5
(a) Serial 54.59 42.26
(b) 54.72 44.74
(c) Double-branch Parallel 54.35 43.33
(d) 55.19 45.01
(e) ✓§ 56.84 45.65
@@
Table 4: Evaluation results of different decoder structures and their Position Encoding methods. §means that the PE method of the decoder’s surrounding branch is Text-guided.
Test of Convergence by Long/Short Epochs Experiments.
To demonstrate the rapid convergence of our model, we conduct ablation experiments with short epochs and exhibit the results in Tab. 5. EDA(Wu et al. 2023) in experiment (a) is our baseline, typically converging around 70 epochs in a 400-epoch experiment. Our PD-TPE can finish convergent around 50 epochs with a higher evaluation result. We also set up 60-epoch experiments for both methods. As shown in experiment (c) and experiment (d), our PD-TPE reaches comparable results as experiment (b), while EDA only achieves 39.72% Acc@0.5IoU. The result reveals that our method can achieve high-precision 3D visual grounding through a short training period.
Epochs Method @0.25 @0.5
(a) 400 EDA 54.59 42.26
(b) PD-TPE 56.84 45.65
(c) 60 EDA † 52.57 39.72
(d) PD-TPE 55.83 45.30
Table 5: Comparision of Convergence between EDA and our PD-TPE. †means that we retrain the model with their published code for 60 epochs.
@@
In Fig. 5, we exhibit the visualization results of 4 scenes from the ScanRefer dataset. The boxes in colors of red, blue, and green represent the ground-truth boxes, the top-1 predicted boxes generated by EDA(Wu et al. 2023), and boxes generated by our PD-TPE. Other light blue boxes are associated with contextual objects according to the text information. The results illustrate the efficacy of our method in comprehending contextual information within the text to infer the referred target object. On the one hand, it can fully perceive the characteristics of the object itself, avoiding errors in extracting valid features such as color and shape (e.g., Fig. 5 (a) identifies the black-and-white monitor in the description as the white one.), as well as errors in object category recognition(e.g., Fig. 5 (d) recognizes ottoman as brench.). On the other hand, it can effectively understand the contextual layout around the target proposals, thereby correctly positioning the described object (e.g., Fig. 5 (b)(c)).
@@ Conclusion
In this paper, we present PD-TPE, a one-stage model with a parallel decoder, for the 3D Visual Grounding task, which helps extract features of proposals and perceive the contextual layouts around them separately. A novel Text-guided Position Encoding method is proposed to guide the generation of the attention map. It is dominated by the relative position between visual tokens and predicted boxes in the main branch, and facilitated by the text tokens with effective layout information in the surrounding branch. Comprehensive experiments and ablation studies demonstrate the effectiveness of our method.

### v2

@@ Introduction
The PE method mentioned above uses relative positions alone to provide confidence for the attention map. This is inaccurate in the surrounding branch, because it cannot determine whether the seed point carries valid surrounding layout information. As shown in Fig. 4, we propose a text-guided method as gates to refine the attention map and update Eq. 1 as follows:
$\widehat{\mathcal{A}}=\operatorname{softmax}({gate}(\mathcal{Q}\mathcal{K}^{\text{T}}{\;+\;\mathcal{E}})).$ (6)
To be specific, we first compute the confidence between seed point features $V\in\mathbb{R}^{N\times D}$ and surrounding text tokens’ features $T_{s}\in\mathbb{R}^{L_{s}\times D}$:
$confidence=V(T_{s})^{\text{T}},$ (7)
where $confidence\in\mathbb{R}^{N\times{L_{s}}}$. Then we take the maximum confidence for each seed point with one of $L_{s}$ text tokens:
$gate_{0}=\sigma(\operatorname{max}(confidence)),$ (8)
where $gate_{0}\in\mathbb{R}^{N}$. $\sigma$ is the sigmoid function, which is employed to map the maximum confidence to a range between 0 and 1. With higher confidence, the gate value approaches 1, allowing the corresponding value in the attention map to be effectively updated by the seed point; conversely, it approaches 0 and the update of the attention map is ignored. Finally, we repeat $gate_{0}$ for $K$ times to generate $gate\in\mathbb{R}^{N\times K}$ and augment the attention map between seed point features and query features.
@@
Following (Wu et al. 2023), we use position and semantic loss for dense alignment, as follows:
$\mathcal{L}_{all}=\mathcal{L}_{pos}+\mathcal{L}_{sem}.$ (9)
Please refer to (Wu et al. 2023) for more details.
@@
We perform experiments on two commonly used datasets, ScanRefer (Chen, Chang, and Nießner 2020) and Nr3D (Achlioptas et al. 2020). Both datasets are derived from ScanNet (Dai et al. 2017), which contains 1201 indoor 3D scenes for training and 312 for validation. ScanRefer provides 51,583 human-annotated language descriptions of 11,046 objects in 800 3D scenes, which is officially divided into a 36,665-sample training set and a 9,508-sample testing set. Nr3D contains 41,503 language annotations, describing 76 object categories from 707 indoor scenes.
For ScanRefer, we adopt Acc@mIoU as the evaluation metric, where m is set as 0.25 and 0.5. This metric indicates the proportion of predicted bounding boxes that have an Intersection over Union (IoU) greater than m with the ground-truth bounding boxes. For Nr3D, experiments are all evaluated under the Acc@0.25IoU.
Ground Truth Protocol Classifier Protocol
Method Easy Hard V-Dep V-Indep Overall Easy Hard V-Dep V-Indep Overall
ReferIt3D (Achlioptas et al. 2020) 43.6 27.9 32.5 37.1 35.6 - - - - -
TGNN (Huang et al. 2021) 44.2 30.6 35.8 38.0 37.3 - - - - -
InstanceRefer (Yuan et al. 2021) 46.0 31.8 34.5 41.9 38.8 - - - - -
3DVG-Trans (Zhao et al. 2021) 48.5 34.8 34.8 43.7 40.8 - - - - -
SAT (Yang et al. 2021) 56.3 42.4 46.9 50.4 49.2 - - - - -
MVT (Huang et al. 2022) 61.3 49.1 54.3 55.4 55.1 - - - - -
BUTD-DETR (Jain et al. 2022) - - - - - 60.7 48.4 46.0 58.0 54.6
ViL3DRel (Wu et al. 2023) 70.2 57.4 62.0 64.5 64.4 - - - - -
EDA (Wu et al. 2023) 58.2 46.1 50.2 53.1 52.1
ViewRefer (Guo et al. 2023) 63.0 49.7 55.1 56.8 56.0 - - - - -
3DRP-Net (Wang et al. 2023) 71.4 59.7 64.2 65.2 65.9 - - - - -
VPP-Net (Shi, Wu, and Lee 2024) - - - - - - - - - 56.9
PD-APE(Ours) 77.0 58.5 64.4 70.1 67.8 67.7 50.6 52.8 62.3 59.1
Table 2: Performance on Nr3D datasets by Acc@0.25IoU as the metric. There are some previous methods that only provide evaluation results on the dataset with detected lables, which are not shown in this table.
PE Method @0.5 Latency GPU Memory
(a) - 42.3 118 ms 3416 MB
(b) Vertex RPE 44.5 155 ms 4168 MB
(c) Center RPE 43.0 121 ms 3668 MB
(d) Box-surface RPE 44.5 123 ms 3672 MB
Table 3: Evaluation results and inference costs of different types of PE methods. For Latency costs per scene and the GPU memory costs, we evaluate on one NVIDIA 4090 24GB GPU with batch size as 1 for a fair comparison.
@@
Tab. 1 exhibits our experiments on the ScanRefer dataset compared with previous works. Our PD-APE with the backbone of PointNet++ outperforms all the previous methods on Acc@0.25IoU and Acc@0.5IoU, achieving 56.6% and 45.5%, respectively. It surpasses our baseline EDA (Wu et al. 2023) by 3.2% Acc@0.5IoU, and also 1.7% higher than that of the previous state-of-the-art method (Yang et al. 2023). Replacing the backbone with Sparse3D ResNet34, our PD-APE reaches even higher results, 57.8% Acc@0.25IoU and 47.1% Acc@0.5IoU. Note that compared to methods using 2D+3D inputs, our approach achieves even better results without relying on additional 2D features for assistance.
We also report our experiment results with PointNet++ on the Nr3d dataset. Different from ScanRefer with detected boxes as ground truth, the grounding task on Nr3d aims to locate target objects from all the given ground truth bounding boxes. In the standard protocol from ReferIt3D (Achlioptas et al. 2020), the ground truth labels for boxes are adopted to independently assess the model’s grounding ability. BUTD-DETR (Jain et al. 2022) uses ground truth boxes with labels from the pointnet++ classifier. For a fair comparison, we provide the evaluation results under both protocols separately in Tab. 2. During the training phase, we employ the data augmentation method in ViL3DRel (Chen et al. 2022b) in the ground truth protocol, and use the data augmentation method following BUTD-DETR (Jain et al. 2022) in the classifier protocol for fairness.
In the ground truth protocol, our PD-APE performs better than all the previous methods and reaches 67.8%, surpassing the previous state-of-the-art method (Wang et al. 2023) by 1.9%. In the classifier protocol, PD-APE achieves an overall accuracy of 59.1%, 2.2% higher than the current best result reported by VPP-Net (Shi, Wu, and Lee 2024) and 7.0% higher than our baseline EDA (Wu et al. 2023).
Decoder
Structure @0.5 Latency/scene GPU Memory
(a) Serial 42.3 118 ms 3416 MB
(b) Parallel 43.3 125 ms 3644 MB
Table 4: Evaluation results of different decoder structures. For Latency costs per scene and the GPU memory costs, we evaluate on one NVIDIA 4090 24GB GPU with batch size as 1 for a fair comparison.
@@
We perform ablation studies to understand the effectiveness of each component of our PD-APE. Without further specification, all experiments are conducted on ScanRefer with 400 epochs, employing PointNet++ as the backbone.
Comparison of Different Relative Positioning Methods.
@@
In Tab. 3, we first compare the effect of different Relative Positioning methods. The experiment (a) is our baseline EDA (Wu et al. 2023). Keeping the structure of the model, we add PE with different types of relative positioning methods to the decoder’s cross-attention block. The experiment (b) use Vertex Relative PE following (Shen et al. 2024), as shown in Fig. 3 (a), achieving 44.5% (+2.2%) Acc@0.5IoU. As shown in experiment (c), the Center Relative PE method (Fig. 3 (b)) results in limited improvement to Acc@0.5IoU (43.0%, +0.7%), because the method only considers the relative positional relationship between seed points and query proposals, without taking the shape and size of the predicted boxes into account. We finally use the Box-surface PE method in experiment (d) (Fig. 3 (c) &(d)), which reaches comparable results as experiment (b). Considering the large latency and unavoidable memory costs brought by Vertex RPE, we take the Box-surface RPE as our priority.
Comparison of Different Structures of Decoders.
@@
We provide the ablation study on the structure of the decoder and illustrate the results in Tab. 4. Comparing the experiment (a) and (b), our design that changes the decoder’s serial structure to a dual-branch parallel structure increases the Acc@0.5IoU from 42.3% to 43.3% (+1.0%). This proves that allowing the decoder to separately focus on the features of the query proposals and their surrounding layouts can lead grounding more accurate.
Figure 5: Visualization results of different models for scenes from the ScanRefer dataset. For all boxes, yellow represents the ground-truth references; red represents results from EDA that contain grounding errors; green represents proposals generated by our PD-APE. Words in different colors show the results of text decoupling.
Effect of Text-guided PE Methods on Two Branches.
To verify the effectiveness of the PE for our dual-branch decoder, we conduct ablation experiments and present the results in Tab. 5. In experiment (a), we adopt the parallel decoder with the naive attention method and no position encoding, which reaches 43.3% Acc@0.5IoU. After equipping Box-surface RPE in both branches in experiment (b), the evaluation result of Acc@0.5IoU achieves 45.0% (+1.7%). Then in experiment (c), we add a text-guided gate to the PE for the surrounding branch in the decoder, which increases the evaluation result to 45.5% Acc@0.5IoU. Note that we also add the text-guided gate in both branches, but the Acc@0.5IoU only reaches 44.8% in experiment (d). The decrease occurs because, for the target object branch, the decoder overlooks points that are incorrectly filtered out by the gate, thus failing to fully learn the query features.
PE for t-branch PE for s-branch @0.25 @0.5
(a) 54.4 43.3
(b) 55.2 45.0
(c) ✓§ 56.6 45.5
(d) ✓§ ✓§ 55.8 44.8
@@
Table 5: Evaluation results of Adaptive Position Encoding methods for different branches. ✓means naive PE method and §means Text-guided PE method.
Effect of Different Text-guided Augmentations.
In this section, we set experiments to evaluate how text-guided confidence can best augment the attention map. In Tab. 6, $attn$ represents the attention map, $PE$ represents the relative position encoding, and $conf$ is the confidence between text tokens and seed points, which is described in detail in Sec. Methodology. Based on experiment (a) with no text-guided confidence, we try to add confidence as bias in experiment (b), and as the gate for PE and the entire attention map in experiment (c) and (d), separately. The results indicate that using confidence as a gate on the entire attention map can best guide queries to attend to seed points that provide effective layout information, achieving 45.5% Acc@0.5IoU.
Text-guided method @0.25 @0.5
(a) - 55.2 45.0
(b) $attn+PE+conf$ 55.8 45.1
(c) $attn+conf\times PE$ 55.6 45.2
(d) $conf\times(attn+PE)$ 56.6 45.5
Table 6: Evaluation results of Different Text-guided method. Experiments above all employ the parallel decoder and only use the Text-guided method in the surrounding branch.
@@
In Fig. 5, we exhibit the visualization results of 4 scenes from the ScanRefer dataset. The boxes in colors of yellow, red, and green represent the ground-truth boxes, the top-1 predicted boxes generated by EDA (Wu et al. 2023), and boxes generated by our PD-APE. The results demonstrate the effectiveness of our method in understanding contextual information in the text to identify the target object. On the one hand, it can avoid errors in extracting features such as colors and shapes (e.g., Fig. 5 (a) identifies the black-and-white monitor as the white one), as well as errors in object category recognition(e.g., Fig. 5 (d) recognizes ottoman as bench). On the other hand, it can effectively understand the spatial layout around the target proposals, thereby correctly positioning the described object (e.g., Fig. 5 (b)(c)).
@@ Conclusion
In this paper, we presented PD-APE, a one-stage decoding framework for the 3D Visual Grounding task, which helps extract features of proposals and perceive the spatial layouts around them separately. A novel Adaptive Position Encoding method was proposed to guide the generation of the attention map. It is dominated by the relative position between seed points and predicted boxes in the target object branch, and facilitated by the text tokens with effective layout information in the surrounding branch. Comprehensive experiments demonstrated the effectiveness of our method.